(a) What is the wavelength (in meters) of a 93.3 MHz Radio wave frequency? (b) What is the frequency (in Hz) of an x-ray beam with wavelength of 0.0100 nm?


1
Expert's answer
2021-07-06T11:38:09-0400

(a) The wavelength of the radio waves broadcast by an FM radio station is

"λ = \frac{c}{f}"

c is the speed of light in vacuum

"c = 3 \times 10^8 \;m/s \\

f= 93.3 \;MHz = 93.3 \times 10^6\;Hz \\

λ = \frac{3 \times 10^8}{93.3 \times 10^6} = 3.21 \;m"

(b)

"f= \frac{c}{λ} \\

λ= 0.0100 \;nm = 0.01 \times 10^{-9}\;m \\

f = \frac{3 \times 10^8}{0.01 \times 10^{-9}} = 300 \times 10^{17} \\

= 3.0 \times 10^{19}\;Hz"